1. If your weight on earth is 1100N, them what would it be on the sun.
2. What’s the angular velocity of the moon in its orbit?
3. A red paint dot on a bicycle tire is free to rotate. What variable represents its angular velocity?
Answer #1

We know,

W=mg

Where W=weight

M=mass

g=acceleration due to gravity

Know,.

1100=m*9.81(g of earth=9.81m/s2)

m=112.25kg

Weight on sun

W=mg

=112.25*274=30755.1N